# Deep Reinforcement Learning for Real-Time Drone Routing in Post-Disaster Road Assessment Without Domain Knowledge

### v1

@@ 6.5 Real-world application
To validate AEDM’s performance on realistic road networks, we conduct experiments using a publicly available transportation network from the transportation networks (https://github.com/bstabler/TransportationNetworks), which represents a large-scale and complex scenario typical in post-disaster assessment. The Anaheim network (416 nodes, 914 links) serves as a metropolitan transportation system, approximating the complexity of city-wide post-disaster road damage assessment, as shown in Figure 12. Following the coordinate normalization procedure described in Section 5.4.1, we convert the geographical coordinates (latitude and longitude) of the network to planar projection coordinates, then apply center translation and equidistant regularization to achieve uniform normalization within $0,1]^{2}$. For Anaheim, we test drone numbers $K\in\{5,6,7\}$ and the maximum allowable assessment time $p_{\max}\in\{30,45,60\}$ minutes with Gurobi’s time limit extended to 10,000 seconds due to increased problem complexity. Figure 13 presents scatter plots comparing AEDM and Gurobi performance across all parameter combinations, with computation time on the x-axis and objective value on the y-axis, demonstrating AEDM’s superiority in both solution quality and computational efficiency. For solution quality, AEDM consistently outperforms Gurobi with gaps of 11.67–24.68% for Anaheim, while for computational efficiency, AEDM achieves real-time performance with approximately 10 seconds, compared to Gurobi’s full allocated time limit of 10,000 seconds without guarantee of optimal solutions. These results confirm AEDM’s practical applicability for time-critical disaster response scenarios, where both solution quality and rapid deployment are essential.
@@ 7 Conclusion
This study addresses the challenge of rapid post-disaster road network damage assessment through drone routing optimization. We propose AEDM that determines high-quality drone assessment routes in real-time without requiring domain-specific algorithmic expertise. The key contributions include: (1) AEDM achieves superior performance compared to commercial solvers, with solution quality improvements of 16–69% while maintaining real-time inference (1–2 seconds vs. 100–2,000 seconds for traditional methods); (2) A simple yet effective network transformation approach that converts a link-based routing problem into an equivalent node-based formulation; (3) A synthetic road network generation method that addresses the scarcity of large-scale training datasets for DRL applications. Experimental results demonstrate AEDM’s robustness across varying problem scales, drone numbers, and time constraints, with consistent generalization to unseen parameter distributions. The model’s ability to balance computational efficiency with solution quality makes it particularly suitable for time-critical disaster response scenarios where rapid decision-making is essential.
Future research can pursue two directions. First, while the proposed multi-task model addresses scenarios with multiple drone numbers and assessment time limits, it can be extended to accommodate more complex road damage assessment scenarios by integrating diverse constraints—including multiple depots, time windows for individual road links, and route “open/closed” assessment status. We plan to develop a unified model to tackle a broader spectrum of road damage assessment problems. Second, drawing on the scaling law in large language models, where performance improves with parameter size (currently reaching billions) (Achiam et al. 2023, Touvron et al. 2023, Guo et al. 2025), we aim to investigate whether increasing the parameter count of our model can further enhance its performance.
Appendices
A. Derivation of point $C$ coordinates
@@
B. Mathematical formulation
To formalize the problem, we define the binary decision variable $x_{ij}^{k}$, where $x_{ij}^{k}=1$ if drone $k$ traverses road link $(i,j)$ for assessment, and 0 otherwise. Using notations and network transformations from Sections 3 and 4, the drone-based road network assessment model is:
@@ B. Mathematical formulation
$\sum\limits_{j\in\bar{N}}{x_{ji}^{k}}=\sum\limits_{j\in\bar{N}}{x_{ij}^{k}}\quad\forall i\in\bar{N},\forall k\in K$ (B3)
$\sum\limits_{j\in\bar{N}}{x_{oj}^{k}}=\sum\limits_{i\in\bar{N}}{x_{io}^{k}}=1\quad\forall k\in K$ (B4)
$\sum\limits_{(i,j)\in\bar{A}}{{t_{ij}}x_{ij}^{k}}\leq{Q}\quad\forall k\in K$ (B5)
$\mathop{\max}\limits_{k\in K}\left\{\sum\limits_{(i,j)\in\bar{A}}{{t_{ij}}x_{ij}^{k}}\right\}\leq p_{max}$ (B6)
$u_{i}^{k}+1-u_{j}^{k}\leq M(1-x_{ij}^{k})\quad\forall(i,j)\in\bar{A},\forall k\in K$ (B7)
$x_{ij}^{k}\in\left\{{0,1}\right\}\quad\forall(i,j)\in\bar{A},\forall k\in K$ (B8)
$u_{i}^{k}\in\left\{{1,2,...,\left|N\right|}\right\}\quad\forall i\in\bar{N},\forall k\in K$ (B9)
  • Equation (B1) maximizes the total information value collected by all drones, summing $c_{p}$ for each transformed artificial node traversed by drone $k$ (as indicated by $x_{ij}^{k}$).
  • Constraints (B2) restrict each information-bearing artificial node to at most one drone visit, preventing redundant assessments by avoiding overlapping visits from multiple drones and thus ensuring assessment efficiency. Notably, this constraint does not impose a single-visit requirement on nodes within the original road network $G$, as illustrated in Figure 3.
@@
  • Constraints (B5) cap drone $k$’s total flight time (sum of $t_{ij}x_{ij}^{k}$) at battery limit $Q$.
@@
In the example in Figure 5, loops are valid and useful. Thus, unlike VRP and TSP, the loop-eliminating Miller-Tucker-Zemlin (MTZ) constraint would ideally be omitted here, as our problem allows loops in paths. However, model inspection reveals a critical issue: solving the model without such constraints yields discontinuous, invalid paths. To resolve this tension—enabling loops while ensuring path continuity—we initially proposed splitting nodes in $N$ into three sub-nodes, with incoming/outgoing edges derived from the original node’s edges. This approach, though, drastically increases node number and complicates solving. Consequently, consultation with domain modeling experts—itself a challenging and time-intensive process—was necessary, leading to a pragmatic decision: this model adopts the MTZ constraint. Specifically, the MTZ constraint in equation (B7) is utilized to preclude cycles or loops in the drone’s path (Miller et al. 1960), enforced by a sufficiently large constant $M$ (typically set to $|\bar{N}|$). Auxiliary variables $u_{i}^{k}$ ensure that if drone $k$ travels from node $i$ to node $j$, the path remains acyclic and continuous, thus avoiding node revisits. Finally, equations (B8) and (B9) define the domains of decision variables: $x_{ij}^{k}\in\{0,1\}$ (binary indicator for whether drone $k$ traverses link $(i,j)$), and $u_{i}^{k}\in\{1,2,...,|\bar{N}|\}$ (integers representing the visit order of nodes by drone $k$).
C. Ablation study
@@
All variants were trained under identical conditions detailed in Section 6.1 and evaluated on 400-node networks with $p_{\text{max}}=45$ minutes and 4, 5, 6, 7 drones (i.e., four scenarios). Performance was quantified by the average optimality gap across these four scenarios, defined as $\text{Gap}=(y-y_{\text{other}})/y$. Here, $y$ denotes the objective value from the EMA with Z-score normalization, and $y_{\text{other}}$ represents that from other normalization variants. As visualized in Figure 14, the proposed reward normalization strategy achieves the best performance, whereas the “No normalization” variant yields the worst. These results underscore the critical role of EMA-based mean and variance tracking in stabilizing reinforcement learning for multi-task drone routing. The proposed normalization strategy ensures reward comparability across diverse parameter regimes (e.g., varying drone numbers and time limits), enabling effective model generalization to unseen scenarios.
@@
D. Method interpretability
To explore the model’s interpretability, we employ t-distributed stochastic neighbor embedding (t-SNE) to project high-dimensional representations—learned under varying drone numbers $k$, $p_{\max}$, and a fixed 100-node scale—into a 2D space. With 100 instances per parameter combination, Figure 15 visualizes the evolution of features across the output of each encoder layer.

### v2

@@ 6.5 Real-world application
To validate AEDM’s performance on realistic road networks, we conduct experiments using a publicly accessible transportation network from the GitHub repository (https://github.com/bstabler/TransportationNetworks), which represents a large-scale and complex scenario typical in post-disaster assessment. The Anaheim network (416 nodes, 914 links) serves as a transportation system, approximating the complexity of post-disaster road damage assessment, as shown in Figure 12. Following the coordinate normalization procedure described in Section 5.4.1, we convert the geographical coordinates (latitude and longitude) of the network to planar projection coordinates, then apply center translation and equidistant regularization to achieve uniform normalization within $0,1]^{2}$.
For Anaheim, we test drone numbers $K\in\{5,6,7\}$ with fixed $p_{\max}=45$ minutes and Gurobi’s time limit extended to 10,000 seconds due to increased problem complexity. Figure 13 presents performance comparisons across all drone configurations. Figure 13 (a) demonstrates solution quality through grouped bar charts, where AEDM×8 consistently achieves the highest objective values (115, 129, 144 for $K=5,6,7$), outperforming Gurobi by 16–27%, GC by 28–32%, and GC+LS by 20–25%. Figure 13 (b) illustrates the efficiency-quality trade-off on logarithmic time scale, demonstrating AEDM’s superiority in both solution quality and computational efficiency. AEDM achieves rapid performance with approximately 10 seconds, compared to Gurobi’s full 10,000-second time limit without guarantee of optimal solutions. These results confirm AEDM’s practical applicability for time-critical disaster response scenarios, where both solution quality and rapid deployment are essential.
@@ 7 Conclusion
This study addresses the challenge of rapid post-disaster road network damage assessment through drone routing optimization. We propose AEDM that determines high-quality drone assessment routes within seconds without requiring domain-specific algorithmic expertise. The key contributions include: (1) AEDM consistently outperforms both commercial solvers and traditional heuristics, achieving 20–71% and 23–35% improvements respectively, while maintaining rapid inference (1–2 seconds vs. 100–2,000 seconds); (2) A simple yet effective network transformation approach that converts a link-based routing problem into an equivalent node-based formulation; (3) A synthetic road network generation method that addresses the scarcity of large-scale training datasets for DRL applications. Experimental results demonstrate AEDM’s robustness across varying problem scales, drone numbers, and time constraints, with consistent generalization to unseen parameter distributions. The model’s ability to balance computational efficiency with solution quality makes it particularly suitable for time-critical disaster response scenarios where rapid decision-making is essential.
Future research can pursue two directions. First, while the proposed multi-task model addresses scenarios with multiple drone numbers and assessment time limits, it can be extended to accommodate more complex road damage assessment scenarios by integrating diverse constraints, including multiple depots, time windows for individual road links, and route “open/closed” assessment status. We plan to develop a unified model to tackle a broader spectrum of road damage assessment problems. Second, inspired by the scaling laws of large language models, which indicate that performance improves as parameter counts grow to the billion scale (achiam2023gpt; touvron2023llama; guo2025deepseek), we aim to investigate whether increasing the parameter count of our model can further enhance its performance.
Appendix A Derivation of point $C$ coordinates
@@
Appendix B Mathematical formulation
To formulate the problem, we define the binary decision variable $x_{ij}^{k}$, where $x_{ij}^{k}=1$ if drone $k$ traverses road link $(i,j)$ for assessment, and 0 otherwise. Additionally, we introduce the continuous flow variable $f_{ij}^{k}\geq 0$ representing the flow carried by drone $k$ on link $(i,j)$. Following the notation and network transformation introduced in Sections 3 and 4, the drone-based road network assessment model is:
@@ Appendix B Mathematical formulation
$\sum_{j\in\bar{N}}x_{ji}^{k}=\sum_{j\in\bar{N}}x_{ij}^{k}\quad\forall i\in\bar{N},\forall k\in K$ (B3)
$\sum_{j\in\bar{N}}x_{oj}^{k}=\sum_{i\in\bar{N}}x_{io}^{k}=1\quad\forall k\in K$ (B4)
$x_{pj}^{k}+x_{jp}^{k}\leq 1\quad\forall j\in\bar{N},\forall p\in\mathcal{P},\forall k\in K$ (B5)
$\sum_{(i,j)\in\bar{A}}t_{ij}x_{ij}^{k}\leq Q\quad\forall k\in K$ (B6)
$\max_{k\in K}\left\{\sum_{(i,j)\in\bar{A}}t_{ij}x_{ij}^{k}\right\}\leq p_{\max}$ (B7)
  • Equation (B1) maximizes the total information value collected by all drones, summing $c_{p}$ for each transformed artificial node traversed by drone $k$.
  • Constraints (B2) restrict each information-bearing artificial node to at most one drone visit, preventing redundant assessments by avoiding overlapping visits from multiple drones. Notably, this constraint does not impose a single-visit requirement on nodes within the original road network $G$, as illustrated in Figure 3.
@@
  • Constraints (B5) forbid a drone from simultaneously using both links $(p,j)$ and $(j,p)$ between an artificial node $p$ and an adjacent node $j$. Since each artificial node $p\in\mathcal{P}$ represents a road link in the original network, a feasible traversal should enter $p$ from one junction and leave towards another junction. Such a pattern is an artifact of the link-to-node transformation. It does not correspond to any meaningful movement in the original road network and is therefore excluded.
  • Constraints (B6) cap drone $k$’s total flight time at battery limit $Q$.
@@
Since the transformed network allows revisiting original nodes while requiring all assessed links to form connected routes from the depot, classical Miller–Tucker–Zemlin (MTZ) subtour elimination constraints used in the existing OP literature (kobeaga2018efficient; zhang2023robust) are unsuitable, because they eliminate all cycles, including those that naturally occur on original nodes (see Figure 5). To simultaneously allow such revisits and prevent disconnected subtours, we propose a single-commodity flow formulation. In this formulation, (i) the depot acts as the unique supply node that provides a total amount of flow equal to the number of artificial nodes visited by a drone; (ii) each visited artificial node consumes exactly one unit of flow, representing the completion of a damage-assessment task; and (iii) original nodes serve as pure transshipment nodes with zero net flow, allowing drones to revisit them without violating flow balance. The following formulation (B8)–(B11) eliminates disconnected subtours while permitting loops at original nodes through the differential treatment of artificial and original nodes in flow conservation constraints, addressing the key modeling challenge that traditional MTZ constraints cannot accommodate.
$\sum_{j:(o,j)\in\bar{A}}f_{oj}^{k}-\sum_{j:(j,o)\in\bar{A}}f_{jo}^{k}=\sum_{p\in\mathcal{P}}\sum_{j:(p,j)\in\bar{A}}x_{pj}^{k}\quad\forall k\in K$ (B8)
$\sum_{j:(j,i)\in\bar{A}}f_{ji}^{k}-\sum_{j:(i,j)\in\bar{A}}f_{ij}^{k}=\sum_{j:(i,j)\in\bar{A}}x_{ij}^{k}\quad\forall i\in\mathcal{P},\forall k\in K$ (B9)
$\sum_{j:(j,i)\in\bar{A}}f_{ji}^{k}-\sum_{j:(i,j)\in\bar{A}}f_{ij}^{k}=0\quad\forall i\in N\setminus\{o\},\forall k\in K$ (B10)
$f_{ij}^{k}\leq|\mathcal{P}|\cdot x_{ij}^{k}\quad\forall(i,j)\in\bar{A},\forall k\in K$ (B11)
$x_{ij}^{k}\in\{0,1\}\quad\forall(i,j)\in\bar{A},\forall k\in K$ (B12)
$f_{ij}^{k}\geq 0\quad\forall(i,j)\in\bar{A},\forall k\in K$ (B13)
  • Constraint (B8) establishes the depot as a supply node, with net outflow equal to the number of artificial nodes visited by drone $k$.
  • Constraint (B9) ensures each visited artificial node consumes one unit of flow, where inflow exceeds outflow by exactly one unit, representing the damage assessment task completion.
  • Constraint (B10) enforces flow conservation at original nodes, where inflow equals outflow, enabling node revisits while maintaining path connectivity. This is the key mechanism that allows drones to revisit original nodes as shown in Figure 5.
  • Constraint (B11) links flow to link usage, ensuring flow is only allowed on links selected by the routing decision.
  • Equations (B12) and (B13) define the domains of decision variables: $x_{ij}^{k}\in\{0,1\}$ is a binary indicator for whether drone $k$ traverses link $(i,j)$, and $f_{ij}^{k}\geq 0$ is the continuous flow variable on link $(i,j)$ for drone $k$.
Appendix C Traditional optimization methods
In Appendix B, we formulated a mathematical model for the drone-based road damage assessment problem, which can be solved using commercial solvers such as Gurobi. To enable a more detailed performance comparison and demonstrate the advantages of the proposed AEDM over traditional optimization methods, we designed a two-phase heuristic algorithm based on established heuristic frameworks for the orienteering problem (OP) (kobeaga2018efficient; yang2025heuragenix). This heuristic consists of a greedy construction phase followed by a local search improvement phase. The construction phase uses a classic profit-to-time ratio rule to generate initial solutions, while the improvement phase applies relocate, exchange, and remove-insert operators that are carefully adapted to the link-to-node transformed network structure. Unlike the proposed AEDM, however, this heuristic requires hand-crafted domain knowledge, illustrating the typical limitations of traditional algorithmic approaches in newly emerging problem settings.
The construction phase builds an initial feasible solution by sequentially assigning routes to each of the $K$ drones. For each drone $k$, starting from the depot $o$, the algorithm iteratively selects the next artificial node $p\in\mathcal{P}\setminus\mathcal{V}$ (where $\mathcal{V}$ denotes the set of already visited artificial nodes) that maximizes the profit-to-time ratio:
$\text{score}(p)=\frac{c_{p}}{\text{ShortestPathTime}(u\to p)}$ (C1)
where $u$ is the current node and $c_{p}$ is the information value of artificial node $p$. The selection is subject to the following constraints: (i) a connectivity constraint requiring that the path $u\to p$ be feasible in the transformed network $\bar{G}$, ensuring that the next node is reachable from the current position; (ii) time feasibility, requiring that the accumulated flight time plus the time to reach $p$ and return to the depot does not exceed the time limit $\min\{Q,p_{\max}\}$; and (iii) exclusivity constraint $p\notin\mathcal{V}$, preventing redundant assessment of the same road link. Once a node is selected, the shortest path from $u$ to $p$ is computed and appended to the current drone’s route. This process continues until no more feasible nodes can be added, at which point the drone returns to the depot and the next drone begins its route construction. The greedy construction phase terminates when all $K$ drones have been assigned routes or all artificial nodes with positive information values have been visited.
After obtaining an initial solution from construction phase, the local search phase attempts to improve solution quality through iterative neighborhood exploration. Given the unique structure of the transformed network, we adapt three local search operators to work with triplet structures. A triplet $(u_{p},p,v_{p})$ represents an artificial node $p$ along with its predecessor $u_{p}$ and successor $v_{p}$ in a drone’s route. The three operators are defined as follows:
  • Relocate: Reverse the direction of a triplet within the same route by changing $(u_{p},p,v_{p})$ to $(v_{p},p,u_{p})$. This operator explores alternative approach directions to the same road link.
  • Remove-insert: Remove a triplet $(u_{p},p,v_{p})$ from its route and insert a different triplet $(u_{q},q,v_{q})$ at a new position. This operator removes an existing assessment task and replaces it with an alternative one, potentially improving the route’s total information value.
  • Exchange: Swap two triplets $(u_{p},p,v_{p})$ and $(u_{q},q,v_{q})$ between two different drone routes, potentially improving the overall solution by reassigning assessment tasks.
The local search procedure iteratively evaluates all possible moves using these three operators. At each iteration, if a feasible neighboring solution with higher total information value is found, the current solution is updated. The search continues for a maximum of $\text{max\_iterations}=1,000$ iterations. The final solution $\Pi^{*}$ with the highest reward $R^{*}$ encountered during the search is returned as the output of the heuristic algorithm.
Algorithm 2 Greedy Heuristic with Local Search
1:Input: Transformed network $\bar{G}=(\bar{N},\bar{A})$, flight times $t_{ij}$, information values $c_{p}$ for $p\in\mathcal{P}$, depot $o$, number of drones $K$, battery flight time limit $Q$, maximum allowable assessment time $p_{\max}$, maximum local-search iterations max_iterationsmax\_iterations
2:Phase 1: Greedy construction of initial solution
3: Build adjacency structure from $\bar{A}$
4: Initialize visited artificial nodes: $\mathcal{V}\leftarrow\emptyset$
5: Initialize solution: $\Pi\leftarrow\emptyset$, total reward $R\leftarrow 0$
6: Initialize global time limit $T_{\max}=\min\{Q,\,p_{\max}\}$
7:for $k=1$ to $K$ do
8:  Construct greedy route $\pi^{k}$:
9:   Set current node $u\leftarrow o$
10:   Iteratively select next artificial node $p\in\mathcal{P}\setminus\mathcal{V}$ maximizing
11:    $\text{score}(p)=\frac{c_{p}}{\text{ShortestPathTime}(u\rightarrow p)}$
12:   Subject to: (i) feasibility of the path $u\to p$ in the transformed network $\bar{G}$, (ii) accumulated time $\leq T_{\max}$, (iii) feasibility of returning to $d$, (iv) $p\notin\mathcal{V}$
13:   Append shortest-path segments to construct complete route $\pi^{k}$
14:   $\mathcal{V}\leftarrow\mathcal{V}\cup\{p\}$
15:  Add $\pi^{k}$ into $\Pi$; update $R$
16:end for
17:Phase 2: Local search improvement
18: Initialize best solution: $\Pi^{*}\leftarrow\Pi$, $R^{*}\leftarrow R$
19:  Define local operators:
20: (a) Relocate: Change a triplet $(u_{p},p,v_{p})$ into $(v_{p},p,u_{p})$ within the same route.
21: (b) Remove-insert: Remove triplet $(u_{p},p,v_{p})$ from its route and insert $(u_{q},q,v_{q})$ at a new position.
22: (c) Exchange: Swap two triplets $(u_{p},p,v_{p})$ and $(u_{q},q,v_{q})$ between routes.
23:for $iteration=1$ to max_iterationsmax\_iterations do
24:  for operator $\in$ {Relocate, Remove-insert, Exchange} do
25:   Generate candidate solution $\Pi^{\prime}$ using operator
26:   if $\Pi^{\prime}$ feasible and $R(\Pi^{\prime})>R(\Pi)$ then
27:    $\Pi\leftarrow\Pi^{\prime}$
28:    if $R(\Pi^{\prime})>R^{*}$ then
29:     $\Pi^{*}\leftarrow\Pi^{\prime}$,   $R^{*}\leftarrow R(\Pi^{\prime})$
30:    end if
31:    break
32:   end if
33:  end for
34:end for
35:Output: Optimized routes $\Pi^{*}$ and reward $R^{*}$
The complete pseudocode for the two-phase heuristic is provided in Algorithm 2. It is important to note that while this heuristic can produce reasonable solutions, its design required substantial domain expertise to: (1) adapt the profit-to-time ratio heuristic to the transformed network structure, (2) develop triplet-based operators that respect the exclusivity constraints of artificial nodes while allowing revisits to original nodes, and (3) carefully tune the search parameters and stopping criteria. This design complexity highlights a key advantage of the proposed AEDM: it eliminates the need for such hand-crafted problem-specific heuristics by learning effective routing policies directly from data.
Appendix D Ablation study
@@
All variants were trained under identical conditions detailed in Section 6.1 and evaluated on 400-node networks with $p_{\text{max}}=45$ minutes and 4, 5, 6, 7 drones (i.e., four scenarios). Performance was quantified by the average relative gap across these four scenarios, defined as $\text{Gap}=(y-y_{\text{other}})/y$. Here, $y$ denotes the objective value from the EMA with Z-score normalization, and $y_{\text{other}}$ represents that from other normalization variants. As visualized in Figure 14, the proposed reward normalization strategy achieves the best performance, whereas the “No normalization” variant yields the worst. These results underscore the critical role of EMA-based mean and variance tracking in stabilizing reinforcement learning for multi-task drone routing. The proposed normalization strategy ensures reward comparability across diverse parameter regimes (e.g., varying drone numbers and time limits), enabling effective model generalization to unseen scenarios.
Appendix E Method interpretability
To explore the model’s interpretability, we employ t-distributed stochastic neighbor embedding (t-SNE) to project high-dimensional representations, which are learned under varying drone numbers $k$, $p_{\max}$, and a fixed 100-node scale, into a 2D space. With 100 instances per parameter combination, Figure 15 visualizes the evolution of features across the output of each encoder layer.